# Supernovae in 2023 (review): breakthroughs by late observations

### v3

@@ 5.2 Core collapse supernovae (CCSNe) in 2023
    According to the JJEM some CCSNRs might have point-symmetric morphologies. This was verified in three CCSNRs in 2023 (Figures 1 - 3). I consider this group of three-point-symmetric CCSNRs (and a speculative identification of a point-symmetric morphology in SNR CTB 1; Bear & Soker 2023) to constitute a breakthrough in our quest for the explosion mechanism of CCSNe, which, I suggest, is the JJEM.
@@ 5.3 Type Ia supernovae (SNe Ia) in 2023
In Section 4 I reviewed the scenarios of normal and peculiar SNe Ia that have been studied in 2023. I presented the classification into six scenarios in Table 1 and criticized studies that even in 2023 still mention only the SD and DD scenarios and ignore other scenarios, in particular the CD and DD-MED scenarios.
@@
    The diversity of both normal and peculiar SNe Ia suggests that more than just two scenarios are involved in forming the majority of all types of normal and peculiar SNe Ia (likely 4 to 5 scenarios). Studies should be open to the different scenarios (Table 1). The habit of mentioning only the SD and DD scenarios should be abandoned.
@@
More research, like the search and detection of supernova progenitors, both CCSNe (e.g., Strotjohann et al. 2023; Van Dyk et al. 2023) and SNe Ia, are essential to verify or reject the breakthroughs I pointed out in this review. Essential to the study of explosion mechanisms and progenitor properties is to consider both theoretical explosion mechanisms of CCSNe and all SN Ia scenarios.
@@ Acknowledgements
I thank Hila Glanz and Amit Kashi for the useful comments. In writing this review I was motivated by the results and discussions of two 2023 meetings on supernovae, SuperNova EXplosions Conference (SNEX, Technion, Israel; chaired by Hagai Perets and Yossef Zenati) and SuperVirtual 2023 (virtual; chaired by Melina Cecilia Bersten and Takashi Moriya). This research was supported by grants from the Pazy Research Foundation and a grant from the Israel Science Foundation (769/20).

### v4

@@ 5.2 Core collapse supernovae (CCSNe) in 2023
    According to the JJEM, some CCSNRs might have point-symmetric morphologies. This was verified in three CCSNRs in 2023 (Figures 1 - 3; for a discussion of the total 8 CCSNRs with identified point-symmetric morphologies, see Soker 2024e). The main caveat in the identification of eight-point-symmetric CCSNRs is the lack of quantitative classification criteria. This is the subject of ongoing research. I consider this group of three-point-symmetric CCSNRs (and a speculative identification of a point-symmetric morphology in SNR CTB 1; Bear & Soker 2023) to constitute a 2023 breakthrough in our quest for the explosion mechanism of CCSNe, which, I suggest, is the JJEM.
@@ 5.3 Type Ia supernovae (SNe Ia) in 2023
In Section 4 I reviewed the scenarios of normal and peculiar SNe Ia that have been studied in 2023. I presented the classification into six scenarios in Table 1 and criticized studies that, even in 2023, still mention only the SD and DD scenarios and ignore other scenarios, particularly the CD and DD-MED scenarios.
@@
    The diversity of both normal and peculiar SNe Ia suggests that more than just two scenarios are involved in forming most of all types of normal and peculiar SNe Ia (likely 4 to 5 scenarios). Studies should be open to the different scenarios (Table 1). The habit of mentioning only the SD and DD scenarios should be abandoned.
@@
More research, like the search and detection of supernova progenitors, both CCSNe (e.g., Strotjohann et al. 2024; Van Dyk et al. 2023) and SNe Ia, are essential to verify or reject the breakthroughs I pointed out in this review. Essential to the study of explosion mechanisms and progenitor properties is to consider both theoretical explosion mechanisms of CCSNe and all SN Ia scenarios.
@@ Acknowledgements
I thank Hila Glanz, Amit Kashi, and two anonymous referees for the useful comments. In writing this review I was motivated by the results and discussions of two 2023 meetings on supernovae, SuperNova EXplosions Conference (SNEX, Technion, Israel; chaired by Hagai Perets and Yossef Zenati) and SuperVirtual 2023 (virtual; chaired by Melina Cecilia Bersten and Takashi Moriya). This research was supported by grants from the Pazy Research Foundation and a grant from the Israel Science Foundation (769/20).